Fermi National Laboratory
 
Operation Summary for April 4, 2012 - April 6, 2012
 

Wednesday April 4

The day shift began with Operations monitoring the Linac, Booster, and Main Injector (MI), and with the NuMI, MiniBooNE and, SeaQuest experiments taking beam.

During the day shift, at 9:58 AM, the NuMI beam permit tripped off on a MI incomplete RF gap, but no RF station had tripped; it reset. At 10:11 AM, the NuMI beam permit tripped off for the first of two times in 30 minutes due to a power supply’s (V118) output drifting low; everything reset. Operations established beam to the Meson FTBFT-992 experiment at 12:19 PM. Operations established beam to MTA. SeaQuest made two accesses during this shift.

During the evening shift, at 4:16 PM, the NuMI beam permit tripped off due to V118; everything reset. Operations disabled the MTA beam at 4:28 PM. Operations disabled the T-992 beam at 6:58 PM. SeaQuest made an access during this shift.

Thursday April 5

The midnight shift began with Operations sending beam to NuMI, MiniBooNE, and SeaQuest.

During the midnight shift, at 5:44 AM, the NuMI beam permit tripped off due to V118; everything reset.

During the day shift, Muon Ring personnel began a study around 8:50 AM. At 11:42 AM, the MiniBooNE beam permit tripped off due to an enclosure air flow indication (E:12AAIR) for the first of two times in about two hours; it reset. Operations established beam to T-992 at 12:11 PM. MI personnel began a slow spill study at 1:21 PM. Operations established beam to MTA around 2 PM. The MI slow spill study ended at 2:38 PM. The Muon Ring study ended at 3:57 PM. SeaQuest and T-992 made accesses during this shift.

During the evening shift, Operations disabled MTA beam at 4:33 PM. At 8:18 PM, Operations reported that the NuMI beam permit had tripped off twice due to V118; it reset. At 8:32 PM, the NuMI beam permit tripped off due to V118’s capacitor bank indication; Operators reported smelling something and called in Power Supply personnel. NuMI resumed taking beam at 10:08 PM; Power Supply technicians reported that they had replaced 10 capacitors and diodes in the V118 power supply. At 10:33 PM, the NuMI beam permit tripped off due to V118 drifting out of tolerance; it reset. T-992 made an access during this shift.

Friday April 6

The midnight shift began with Operations sending beam to NuMI, MiniBooNE, SeaQuest, and T-992.

During the midnight shift, Operations disabled T-992 beam at 1:24 AM. At 3:58 AM, the NuMI beam permit tripped off for the first of two times in three hours due to V118; it reset.

Operations Coordinator (OpsCo) Plans for Today and the Weekend

The OpsCo agrees with Booster, MI, and Muon Ring study plans for today.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 04/06/2012    email Fermilab
Security, Privacy, Legal  |  Use of CookiesFermi National Accelerator Laboratory